# Sub-optimality of the Separation Principle for Quadratic Control from Bilinear Observations

### v1

@@ 2.1 Bilinear Observation Optimal Control Problem
where ${\bm{x}}_{t}\in\mathbb{R}^{n}$ is the state, ${\bm{u}}_{t}\in\mathbb{R}^{p}$ is the input, ${\bm{y}}_{t}\in\mathbb{R}^{m}$ is the output, ${\bm{w}}_{t}\in\mathbb{R}^{n}$ is the process noise, and ${\bm{z}}_{t}\in\mathbb{R}^{m}$ is the measurement noise at time $t$. The matrices ${\bm{A}}\in\mathbb{R}^{n\times n}$, and ${\bm{B}}\in\mathbb{R}^{n\times p}$ govern the state evolution, whereas, ${\bm{C}}_{0},{\bm{C}}_{1},\dots,{\bm{C}}_{p}\in\mathbb{R}^{m\times n}$ govern the bilinear observation model. For notational convenience, we define the input dependent observation matrix ${\bm{C}}({\bm{u}}_{t}):={\bm{C}}_{0}+\sum_{k=1}^{p}({\bm{u}}_{t})_{k}{\bm{C}}_{k}$. In the remainder of the paper, we use LDS-BO$({\bm{A}},{\bm{B}},{\bm{C}}_{0},{\bm{C}}_{1},\dots,{\bm{C}}_{p})$ to refer to the state-space representation (2.1). With these definitions, we want to analyze the following finite-horizon optimal control problem:
@@
[…]
@@ 2.2 Input-Dependent State Estimation
{mdframed}
Bilinear Observation Kalman Filtering: Let $\hat{{\bm{x}}}_{t|t-1}:=\operatorname{\mathbb{E}}[{\bm{x}}_{t}|\mathcal{I}_{t}]$ be the estimated state at time $t$, and ${\bm{\Sigma}}_{t|t-1}:=\operatorname{\mathbb{E}}\left[({\bm{x}}_{t}-\operatorname{\mathbb{E}}[{\bm{x}}_{t}|\mathcal{I}_{t}])({\bm{x}}_{t}-\operatorname{\mathbb{E}}[{\bm{x}}_{t}|\mathcal{I}_{t}])^{\top}\right]$ be the estimation error covariance. Then, starting from $\hat{{\bm{x}}}_{0|-1}=\hat{{\bm{x}}}_{0}$, and ${\bm{\Sigma}}_{0|-1}={\bm{\Sigma}}_{0}$, the estimated state, and the error covariance follows the recursion:
$\hat{{\bm{x}}}_{t+1|t}$ $={\bm{A}}\hat{{\bm{x}}}_{t|t-1}+{\bm{B}}{\bm{u}}_{t}-{\bm{L}}({\bm{u}}_{t})\left({\bm{y}}_{t}-{\bm{C}}({\bm{u}}_{t})\hat{{\bm{x}}}_{t|t-1}\right),$
${\bm{\Sigma}}_{t+1|t}$ $={\bm{A}}{\bm{\Sigma}}_{t|t-1}{\bm{A}}^{\top}+{\bm{L}}({\bm{u}}_{t}){\bm{C}}({\bm{u}}_{t}){\bm{\Sigma}}_{t|t-1}{\bm{A}}^{\top}+{\bm{\Sigma}}_{w},$
[…]
is the input-dependent Kalman gain at time $t$. Please refer to [4, Eqs. (E.39) – (E.42)] for the derivation of Kalman filtering equations above. Importantly, unlike Kalman filtering from linear measurements, both the Kalman gain ${\bm{L}}({\bm{u}}_{t})$, and the error covariance ${\bm{\Sigma}}_{t+1|t}$ depend on the control inputs up to time $t$. Hence, the separation principle (Def. 2) is not valid when controlling the partially observed LDS from bilinear observations (2.1). However, it is easy to show that, the Kalman filtering gives the optimal state estimation of (2.1), under Assumption 1.
@@ 3 Main Results
For the sake of completeness, in the following we present the sub-optimal LQG controller [4] for the system (2.1): at every time $10\leq t\leq T-$ 1, the LQG controller is given by the sub-optimal policy:
@@ 3.1 Sub-optimality of the Separation Principle
${\bm{u}}_{T-2}^{\star}=\arg\min_{{\bm{u}}_{T-2}}f({\bm{u}}_{T-2})$ $:=f_{\rm LQG}({\bm{u}}_{T-2})+g({\bm{u}}_{T-2}),$ (3.3)
$\text{where}\quad f_{\rm LQG}({\bm{u}}_{T-2})$ $:={\bm{u}}_{T-2}^{\top}\mathcal{A}{\bm{u}}_{T-2}+2\hat{{\bm{x}}}_{T-2|T-3}^{\top}\mathcal{B}^{\top}{\bm{u}}_{T-2},$
$\text{and}\;\;\quad\quad\quad g({\bm{u}}_{T-2})$ $:=\textup{{tr}}\big{(}\mathcal{G}\big{(}{\bm{\Sigma}}_{T-2|T-3}^{-1}+{\bm{C}}({\bm{u}}_{T-2})^{\top}{\bm{\Sigma}}_{z}^{-1}{\bm{C}}({\bm{u}}_{T-2})\big{)}^{-1}\big{)},$
where we define $\mathcal{A}:={\bm{B}}^{\top}{\bm{K}}_{T-1}{\bm{B}}+{\bm{R}}$, $\mathcal{B}:={\bm{B}}^{\top}{\bm{K}}_{T-1}{\bm{A}}$, and $\mathcal{G}:={\bm{A}}^{\top}{\bm{P}}_{T-1}{\bm{A}}$ for notational convenience. Note that the second term $g({\bm{u}}_{T-2})$ depends on control input ${\bm{u}}_{T-2}$ via the input dependent observation matrix ${\bm{C}}({\bm{u}}_{T-2})$, and it corresponds to the estimation error covariance at time $T-1$. Hence, the control input at time $T-2$ directly affects the estimation error covariance at time $T-1$, violating the separation principle. Our next result shows that, because of the additional term $g({\bm{u}}_{T-2})$, the optimal control policy is not affine in the estimated state.
@@ 4.1 Proof of Theorem 1
=$\stackrel{{\scriptstyle\text{(i)}}}{{=}}{\operatorname{\mathbb{E}}}\left[\left(x_{T-1}-\hat{x}_{T-1|T-2}\right)^{2}\big{|}\mathcal{I}_{T-2},u_{T-2}\right],$
@@
=$\stackrel{{\scriptstyle\text{(ii)}}}{{=}}A^{2}\Sigma_{T-2|T-3}-\frac{A^{2}(C_{0}+C_{1}u_{T-2})^{2}\Sigma_{T-2|T-3}^{2}}{(C_{0}+C_{1}u_{T-2})^{2}\Sigma_{T-2|T-3}+\Sigma_{z}}+\Sigma_{w},$
@@
$u_{T-2}^{\star}=\arg\min_{u}f_{T-2}(u):=\alpha u^{2}+2\beta\hat{x}u+\frac{\gamma}{(C_{0}+C_{1}u)^{2}+\kappa},$ (4.5)
@@
$\alpha$ $:=B^{2}K_{T-1}+R\qquad\qquad\beta:=BK_{T-1}A$ (4.6)
$\gamma$ $:=\Sigma_{z}A^{2}P_{T-1}\qquad\qquad~{}~{}\kappa:=\Sigma_{z}\Sigma_{T-2|T-3}^{-1}$
@@
$\bar{u}_{T-2}^{\star}=\arg\min_{\bar{u}}\bar{f}_{T-2}(\bar{u})$ $:=\alpha\left(\frac{\bar{u}-C_{0}}{C_{1}}\right)^{2}+2\beta\hat{x}\left(\frac{\bar{u}-C_{0}}{C_{1}}\right)+\frac{\gamma}{\bar{u}^{2}+\kappa},$ (4.7)
@@
=$\stackrel{{\scriptstyle\text{(i)}}}{{=}}(\alpha\bar{L}+C_{1}\beta)C_{0}^{4}+2\kappa(\alpha\bar{L}+C_{1}\beta)C_{0}^{2}+\kappa^{2}(\alpha\bar{L}+C_{1}\beta)-\gamma C_{1}^{2}\bar{L},$
=$\stackrel{{\scriptstyle\text{(ii)}}}{{=}}\frac{\gamma C_{1}^{3}\beta}{\alpha}\neq 0,$ (4.13)
@@ 4.2 Proof of Theorem 2
Using the assumption that $-\frac{C_{0}}{C_{1}}=-\frac{\beta\hat{x}_{0}}{\alpha}=:u_{0}^{LQG}$, we get $C_{1}\beta\hat{x}_{0}-C_{0}\alpha=0$. This further implies that the optimal $\bar{u}_{0}^{\star}$ is one of the roots of the following polynomial,
@@
$\alpha\bar{u}^{4}+2\alpha\kappa\bar{u}^{2}+\alpha\kappa^{2}-\gamma C_{1}^{2}=0\implies\bar{u}_{0}^{\star}$ =±$=\pm\sqrt{-\kappa\pm C_{1}\sqrt{\frac{\gamma}{\alpha}}},$
@@ 4.3 Proof of Proposition 1
To begin, we use [4, Eq. (4.4)] to find the optimal control input at time $T-1$ as follows,
@@
=$\stackrel{{\scriptstyle\text{(i)}}}{{=}}{\bm{A}}{\bm{\Sigma}}_{T-2|T-3}^{1/2}\left({\bm{I}}_{n}+{\bm{\Sigma}}_{T-2|T-3}^{1/2}{\bm{C}}({\bm{u}}_{T-2})^{\top}{\bm{\Sigma}}_{z}^{-1}{\bm{C}}({\bm{u}}_{T-2}){\bm{\Sigma}}_{T-2|T-3}^{1/2}\right)^{-1}{\bm{\Sigma}}_{T-2|T-3}^{1/2}{\bm{A}}^{\top}+{\bm{\Sigma}}_{w},$
@@
Finally, using (4.29) in (4.26), and after some algebraic manipulations, the optimal control input at time $T-2$ is given by the following non-convex minimization problem,
[…]
@@ 5 Numerical Experiments
 (a) $-\frac{C_{0}}{C_{1}}=-\frac{\beta\hat{x}_{0}}{\alpha}+0.5$  (b) $-\frac{C_{0}}{C_{1}}=-\frac{\beta\hat{x}_{0}}{\alpha}$  (c) $-\frac{C_{0}}{C_{1}}=-\frac{\beta\hat{x}_{0}}{\alpha}-0.5$  (d) $-\frac{C_{0}}{C_{1}}=-\frac{\beta\hat{x}_{0}}{\alpha}-1.0$
@@ 5.1.1 Scalar Example
[…]
@@ 5.1.2 Double Integrator Style Dynamics
[…]
@@ 5.2 Scalar Example
Figure 1 plots $f_{0}(u)$ defined in (4.14) over different values of $u$. Our plots show that the critical points of $f_{0}(u)$ depends on the distance between the optimal LQG controller $u_{0}^{\rm LQG}=-\frac{\beta\hat{x}_{0}}{\alpha}$ and the global maxima of the estimation error variance term $\frac{\gamma}{(C_{0}+C_{1}u)^{2}+\kappa}$, which is $-\frac{C_{0}}{C_{1}}$ in this example. Interestingly, Figure 1 shows that when $\frac{C_{0}}{C_{1}}=\frac{\beta\hat{x}_{0}}{\alpha}$, the LQG controller $u_{0}^{\rm LQG}=-\frac{\beta\hat{x}_{0}}{\alpha}$ locally maximizes $f_{0}(u)$, which is non-convex, However, when $\frac{C_{0}}{C_{1}}$ is far from $\frac{\beta\hat{x}_{0}}{\alpha}$, the optimal nonlinear controller converges to the LQG controller.

### v2

@@ 2.1 Bilinear Observation Optimal Control Problem
[…]
@@
We have: (i) Gaussian noise processes, $\{{\bm{w}}_{t}\}_{t=0}^{T-1}\overset{\text{i.i.d.}}{\sim}\mathcal{N}(0,{\bm{\Sigma}}_{w})$, $\{{\bm{z}}_{t}\}_{t=0}^{T-1}\overset{\text{i.i.d.}}{\sim}\mathcal{N}(0,{\bm{\Sigma}}_{z})$, (ii) initial state distributed as ${\bm{x}}_{0}\overset{\text{i.i.d.}}{\sim}\mathcal{N}(\hat{{\bm{x}}}_{0},{\bm{\Sigma}}_{0})$, and (iii) the positive definite cost matrices, i.e., ${\bm{Q}}_{T},{\bm{Q}},{\bm{R}}\succ 0$.
@@ 2.2 Input-Dependent State Estimation
Bilinear Observation Kalman Filtering: Let $\hat{{\bm{x}}}_{t|t-1}:=\operatorname{\mathbb{E}}[{\bm{x}}_{t}|\mathcal{I}_{t}]$ be the estimated state at time $t$, and ${\bm{\Sigma}}_{t|t-1}:=\operatorname{\mathbb{E}}\left[({\bm{x}}_{t}-\operatorname{\mathbb{E}}[{\bm{x}}_{t}|\mathcal{I}_{t}])({\bm{x}}_{t}-\operatorname{\mathbb{E}}[{\bm{x}}_{t}|\mathcal{I}_{t}])^{\top}\right]$ be the estimation error covariance. Then, starting from $\hat{{\bm{x}}}_{0|-1}=\hat{{\bm{x}}}_{0}$, and ${\bm{\Sigma}}_{0|-1}={\bm{\Sigma}}_{0}$, the estimated state, and the error covariance follows the recursion: $\hat{{\bm{x}}}_{t+1|t}$ $={\bm{A}}\hat{{\bm{x}}}_{t|t-1}+{\bm{B}}{\bm{u}}_{t}-{\bm{L}}({\bm{u}}_{t})\left({\bm{y}}_{t}-{\bm{C}}({\bm{u}}_{t})\hat{{\bm{x}}}_{t|t-1}\right),$ ${\bm{\Sigma}}_{t+1|t}$ $={\bm{A}}{\bm{\Sigma}}_{t|t-1}{\bm{A}}^{\top}+{\bm{L}}({\bm{u}}_{t}){\bm{C}}({\bm{u}}_{t}){\bm{\Sigma}}_{t|t-1}{\bm{A}}^{\top}+{\bm{\Sigma}}_{w},$ $\text{where,}\quad{\bm{L}}({\bm{u}}_{t})$ $=-{\bm{A}}{\bm{\Sigma}}_{t|t-1}{\bm{C}}({\bm{u}}_{t})^{\top}\left({\bm{C}}({\bm{u}}_{t}){\bm{\Sigma}}_{t|t-1}{\bm{C}}({\bm{u}}_{t})^{\top}+{\bm{\Sigma}}_{z}\right)^{-1}$ is the input-dependent Kalman gain at time $t$.
Please refer to [4, Eqs. (E.39) – (E.42)] for the derivation of Kalman filtering equations above. Importantly, unlike Kalman filtering from linear measurements, both the Kalman gain ${\bm{L}}({\bm{u}}_{t})$, and the error covariance ${\bm{\Sigma}}_{t+1|t}$ depend on the control inputs up to time $t$. Hence, the separation principle (Def. 2) is not valid when controlling the partially observed LDS from bilinear observations (2.1). However, it is easy to show that, the Kalman filtering gives the optimal state estimation of (2.1), under Assumption 1.
@@ 3 Main Results
For the sake of completeness, in the following we present the sub-optimal LQG controller [4] for the system (2.1): at every time $0\leq t\leq T-1$, the LQG controller is given by the sub-optimal policy:
@@ 3.1 Sub-optimality of the Separation Principle
${\bm{u}}_{T-2}^{\star}=\arg\min_{{\bm{u}}_{T-2}}f({\bm{u}}_{T-2})$ $=f_{\rm LQG}({\bm{u}}_{T-2})+g({\bm{u}}_{T-2}),$ (3.3)
$\text{where}\quad f_{\rm LQG}({\bm{u}}_{T-2})$ $={\bm{u}}_{T-2}^{\top}\mathcal{A}{\bm{u}}_{T-2}+2\hat{{\bm{x}}}_{T-2|T-3}^{\top}\mathcal{B}^{\top}{\bm{u}}_{T-2},$
$\text{and}\;\;\quad\quad\quad g({\bm{u}}_{T-2})$ $=\textup{{tr}}\big(\mathcal{G}\big({\bm{\Sigma}}_{T-2|T-3}^{-1}+{\bm{C}}({\bm{u}}_{T-2})^{\top}{\bm{\Sigma}}_{z}^{-1}{\bm{C}}({\bm{u}}_{T-2})\big)^{-1}\big),$
where we define $\mathcal{A}:={\bm{B}}^{\top}{\bm{K}}_{T-1}{\bm{B}}+{\bm{R}}$, $\mathcal{B}:={\bm{B}}^{\top}{\bm{K}}_{T-1}{\bm{A}}$, and $\mathcal{G}:={\bm{A}}^{\top}{\bm{P}}_{T-1}{\bm{A}}$ for notational convenience. Note that the second term $g({\bm{u}}_{T-2})$ depends on control input ${\bm{u}}_{T-2}$ via the input dependent observation matrix ${\bm{C}}({\bm{u}}_{T-2})$, and it corresponds to the estimation error covariance at time T1T-1. Hence, the control input at time T2T-2 directly affects the estimation error covariance at time T1T-1, violating the separation principle. Our next result shows that, because of the additional term $g({\bm{u}}_{T-2})$, the optimal control policy is not affine in the estimated state.
@@ 4.1 Proof of Theorem 1
$\stackrel{{\scriptstyle\text{(i)}}}{{=}}{\operatorname{\mathbb{E}}}\left[\left(x_{T-1}-\hat{x}_{T-1|T-2}\right)^{2}\big|\mathcal{I}_{T-2},u_{T-2}\right],$
@@
$\stackrel{{\scriptstyle\text{(ii)}}}{{=}}A^{2}\Sigma_{T-2|T-3}-\frac{A^{2}(C_{0}+C_{1}u_{T-2})^{2}\Sigma_{T-2|T-3}^{2}}{(C_{0}+C_{1}u_{T-2})^{2}\Sigma_{T-2|T-3}+\Sigma_{z}}+\Sigma_{w},$
@@
$u_{T-2}^{\star}=\arg\min_{u}f_{T-2}(u)=\alpha u^{2}+2\beta\hat{x}u+\frac{\gamma}{(C_{0}+C_{1}u)^{2}+\kappa},$ (4.5)
@@
$\alpha$ $=B^{2}K_{T-1}+R\qquad\qquad\beta=BK_{T-1}A$ (4.6)
$\gamma$ $=\Sigma_{z}A^{2}P_{T-1}\qquad\qquad~~\kappa=\Sigma_{z}\Sigma_{T-2|T-3}^{-1}$
@@
$\bar{u}_{T-2}^{\star}=\arg\min_{\bar{u}}\bar{f}_{T-2}(\bar{u})$ $=\alpha\left(\frac{\bar{u}-C_{0}}{C_{1}}\right)^{2}+2\beta\hat{x}\left(\frac{\bar{u}-C_{0}}{C_{1}}\right)+\frac{\gamma}{\bar{u}^{2}+\kappa},$ (4.7)
@@
$\stackrel{{\scriptstyle\text{(i)}}}{{=}}(\alpha\bar{L}+C_{1}\beta)C_{0}^{4}+2\kappa(\alpha\bar{L}+C_{1}\beta)C_{0}^{2}+\kappa^{2}(\alpha\bar{L}+C_{1}\beta)-\gamma C_{1}^{2}\bar{L},$
$\stackrel{{\scriptstyle\text{(ii)}}}{{=}}\frac{\gamma C_{1}^{3}\beta}{\alpha}\neq 0,$ (4.13)
@@ 4.2 Proof of Theorem 2
Using the assumption that $\frac{C_{0}}{C_{1}}=-\frac{\beta\hat{x}_{0}}{\alpha}=:u_{0}^{LQG}$, we get $C_{1}\beta\hat{x}_{0}-C_{0}\alpha=0$. This further implies that the optimal $\bar{u}_{0}^{\star}$ is one of the roots of the following polynomial,
@@
$\alpha\bar{u}^{4}+2\alpha\kappa\bar{u}^{2}+\alpha\kappa^{2}-\gamma C_{1}^{2}=0\implies\bar{u}_{0}^{\star}$ $=\pm\sqrt{-\kappa\pm C_{1}\sqrt{\frac{\gamma}{\alpha}}},$
@@ 4.3 Proof of Proposition 1
To begin, we use [4, Eq. (4.4)] to find the optimal control input at time T1T-1 as follows,
@@
$\stackrel{{\scriptstyle\text{(i)}}}{{=}}{\bm{A}}{\bm{\Sigma}}_{T-2|T-3}^{1/2}\left({\bm{I}}_{n}+{\bm{\Sigma}}_{T-2|T-3}^{1/2}{\bm{C}}({\bm{u}}_{T-2})^{\top}{\bm{\Sigma}}_{z}^{-1}{\bm{C}}({\bm{u}}_{T-2}){\bm{\Sigma}}_{T-2|T-3}^{1/2}\right)^{-1}{\bm{\Sigma}}_{T-2|T-3}^{1/2}{\bm{A}}^{\top}+{\bm{\Sigma}}_{w},$
@@
Finally, using (4.29) in (4.26), and after some algebraic manipulations, the optimal control input at time T2T-2 is given by the following non-convex minimization problem,
${\bm{u}}_{T-2}^{\star}=\arg\min_{{\bm{u}}_{T-2}}\bigg\{$ ${\bm{u}}_{T-2}^{\top}{\cal{A}}{\bm{u}}_{T-2}+2\hat{{\bm{x}}}_{T-2|T-3}^{\top}\mathcal{B}^{\top}{\bm{u}}_{T-2}$
@@ 5 Numerical Experiments
   (a) $\frac{C_{0}}{C_{1}}=-\frac{\beta\hat{x}_{0}}{\alpha}+0.5$    (b) $\frac{C_{0}}{C_{1}}=-\frac{\beta\hat{x}_{0}}{\alpha}$    (c) $\frac{C_{0}}{C_{1}}=-\frac{\beta\hat{x}_{0}}{\alpha}-0.5$    (d) $\frac{C_{0}}{C_{1}}=-\frac{\beta\hat{x}_{0}}{\alpha}-1.0$
@@ 5.1.1 Scalar Example
In this experiment, we consider an instance of (2.1) with $n=m=p=1$, and $T=2$. Specifically, we set $A=0.9$, $B=1$, $Q_{T}=Q=R=1$. The initial state is sampled from $x_{0}{\sim}\mathcal{N}(0.1,2)$, whereas, $\{w_{0},w_{1}\}{\overset{\text{i.i.d.}}{\sim}}\mathcal{N}(0,0.01)$, and $\{z_{0},z_{1}\}{\overset{\text{i.i.d.}}{\sim}}\mathcal{N}(0,0.09)$. This gives, $\alpha=2.405$, $\beta=0.126$, $\gamma=0.03$, and $\kappa=0.045$ in (4.6). $C_{1}$ is fixed to $C_{1}=\alpha$, whereas, $C_{0}$ is varied between (3,3)(-3,3) such that $|\frac{C_{0}}{C_{1}}-\frac{\beta\hat{x}_{0}}{\alpha}|\in[0,1]$.
@@ 5.1.2 Double Integrator Style Dynamics
where ${\bm{w}}_{t}{\overset{\text{i.i.d.}}{\sim}}\mathcal{N}(0,0.01{\bm{I}}_{2})$, $z_{t}{\overset{\text{i.i.d.}}{\sim}}\mathcal{N}(0,0.01)$, and we use $h=0.3$. For quadratic cost, we use ${\bm{Q}}={\bm{Q}}_{T}={\bm{I}}_{2}$, and ${\bm{R}}=1000$. The initial state distribution is $\mathcal{N}(0,{\bm{I}}_{2})$. These dynamics approximate a double integrator with discretization step $h$. For example, if the state contains position and velocity, then the input is a force. The bilinear observation model posits that the signal-to-noise ratio of the position sensor scales with the input force. Put differently, the position measurement results in the application of a force.
@@ 5.2 Scalar Example
Figure 1 plots $f_{0}(u)$ defined in (4.14) over different values of $u$. Our plots show that the critical points of $f_{0}(u)$ depends on the distance between the optimal LQG controller $u_{0}^{\rm LQG}=-\frac{\beta\hat{x}_{0}}{\alpha}$ and the global maxima of the estimation error variance term $\frac{\gamma}{(C_{0}+C_{1}u)^{2}+\kappa}$, which is $\frac{C_{0}}{C_{1}}$ in this example. Interestingly, Figure 1 shows that when $\frac{C_{0}}{C_{1}}=\frac{\beta\hat{x}_{0}}{\alpha}$, the LQG controller $u_{0}^{\rm LQG}=-\frac{\beta\hat{x}_{0}}{\alpha}$ locally maximizes $f_{0}(u)$, which is non-convex, However, when $\frac{C_{0}}{C_{1}}$ is far from $\frac{\beta\hat{x}_{0}}{\alpha}$, the optimal nonlinear controller converges to the LQG controller.